OpenStudy (tumblewolf):

I need some help A student answers all 48 questions on a multiple-choice test by guessing. Each question has four possible answers, only one of which is correct. Find the probability that the student gets exactly 15 correct answers. Use the normal distribution to aproximate the binomial distribution. I used BinomCdf(48,.25,15) and get 0.0767 but that isn't one of the choices. The choices are 0.7967 0.0606 0.0823 0.8577

jimthompson5910 (jim_thompson5910):

you use the binomPdf function since you want exactly one probability. You aren't summing a bunch of binomial probabilities

jimthompson5910 (jim_thompson5910):

if it asked something like "what is the probability the student gets at most 15 correct", then you'd be computing P(X <= 15) which would lead you to using binomCdf

OpenStudy (tumblewolf):

I'm sorry I was using PDF, CDF got me 0.8768. I accidentally typed it in wrong

jimthompson5910 (jim_thompson5910):

oh and it says "Use the normal distribution to aproximate the binomial distribution. " I missed that part

jimthompson5910 (jim_thompson5910):

mean = n*p = 48*0.25 = 12 standard deviation = sqrt(n*p*(1-p)) = sqrt(48*0.25*(1-0.25)) = 3

jimthompson5910 (jim_thompson5910):

so you'll use the normal distribution with mean mu = 12 and std dev sigma = 3 to approximate this binomial distribution

OpenStudy (tumblewolf):

So I would use normalpdf?

OpenStudy (tumblewolf):

That's getting me 0.0807

jimthompson5910 (jim_thompson5910):

normal cdf actually look at this page https://people.richland.edu/james/lecture/m170/ch07-bin.html don't forget about the `Continuity Correction Factor`

OpenStudy (tumblewolf):

What would I make my boundaries? I know about the add.5 and subtract .5, I'm not given much to work with

jimthompson5910 (jim_thompson5910):

you'd have the lower boundary be 14.5 the upper boundary would be 15.5

jimthompson5910 (jim_thompson5910):

I'm getting 0.0807 when I compute the area under the curve from 14.5 to 15.5

OpenStudy (tumblewolf):

That's what I'm getting also. It's not a choice though :-\ see where I'm getting confused?

jimthompson5910 (jim_thompson5910):

the more accurate value of P(X = 15) is 0.0767 I'm using the geogebra web app (the probability subsection) to compute these probabilities https://web.geogebra.org/#probability

jimthompson5910 (jim_thompson5910):

strange how none of the listed choice match

OpenStudy (tumblewolf):

That's really what threw me off. I've tried just about everything and nothing is matching out.

jimthompson5910 (jim_thompson5910):

maybe they want you to use a table instead of a calc?

OpenStudy (tumblewolf):

No our teacher hates the table, any time they ask us to use it she makes us use the calculator. We didn't really learn how to do anything with the tables.

jimthompson5910 (jim_thompson5910):

hmm let me try something x = 14.5 z = (x-mu)/sigma z = (14.5-12)/3 z = 0.83333 z = 0.83 x = 15.5 z = (x-mu)/sigma z = (15.5-12)/3 z = 1.1666 z = 1.17 use a table like this one http://www.stat.ufl.edu/~athienit/Tables/Ztable.pdf to find that P(z < 0.83) = 0.7967 P(z < 1.17) = 0.9525 So P(0.83 < z < 1.17) = P(z < 1.17) - P(z < 0.83) P(0.83 < z < 1.17) = 0.8790 - 0.7967 P(0.83 < z < 1.17) = 0.0823

jimthompson5910 (jim_thompson5910):

that's odd how I was able to land on 0.0823 but your teacher didn't show you how to use the table or that she hates the tables. I don't blame her. I prefer the calculator too

jimthompson5910 (jim_thompson5910):

sorry I made a typo, I meant to say P(z < 1.17) = 0.8790

OpenStudy (tumblewolf):

Maybe she's not the one who wrote out the questions? Idk but Thank you so much!!!!! I've been stressing over this for hours.

OpenStudy (thebossofme191):

hours!

jimthompson5910 (jim_thompson5910):

That happens a lot. Often the teacher uses questions found online somewhere or they were written by the department chair (those questions are then shared amongst the whole department to make things easier)

OpenStudy (tumblewolf):

well I went back and finished all the others. We were assigned 400 problems to. (I would hope she didn't write all of them)

OpenStudy (thebossofme191):

the ppl who make tests can kiss butt